# Benchmarking Foundation Models and Parameter-Efficient Fine-Tuning for Prognosis Prediction in Medical Imaging

### v1

@@ 3 Materials & Methods
We employ four publicly available COVID-19 CXR datasets annotated with a variety of prognostic outcomes, including mortality, severity, and ICU admission. These datasets were selected to cover a wide range of clinical scenarios and data challenges, including varying degrees of class imbalance, outcome heterogeneity, and differences in sample size—ranging from highly underrepresented cases to larger, more balanced datasets. This diversity reflects the complexities typically encountered in real-world prognosis modeling and enables a robust evaluation of model adaptability under different data regimes [63, 64]. Further details of the datasets are provided in Section 3.1.
@@
We systematically evaluate models grouped into two main architectural families. The first includes standard CNNs pretrained on ImageNet using supervised learning, representing strong baselines in medical imaging due to their inductive biases and efficiency in low-data scenarios. The second architectural category, FMs), is composed by models pretrained through self-supervised or contrastive learning, either on large-scale general image-text datasets or specialized medical datasets. This categorization enables us to explore the interplay between architectural design, pretraining paradigms, and domain specificity in prognosis prediction tasks.
@@
Each model undergoes adaptation to the target prognostic tasks using one of several fine-tuning strategies. These include FFT as a performance upper bound, LP to evaluate the quality of frozen representations, and multiple PEFT techniques, which are designed to minimize computational overhead and improve adaptation efficiency. The full range of supported configurations and their applicability to specific architectures are detailed in Section 3.
@@
To assess the generalization capabilities of models under realistic constraints, we benchmark each configuration under two training regimes. In the full-data regime, models are trained using all available labeled samples. In contrast, the FSL regime limits training to as few as 2, 4, 8, or 16 samples per class, simulating extreme data scarcity common in rare disease contexts or emerging health crises. This dual evaluation framework enables a comprehensive analysis of model robustness and the effectiveness of fine-tuning strategies across a spectrum of data availability scenarios.
@@
We evaluate model performance primarily using the Matthews Correlation Coefficient (MCC), a metric known for its robustness in the presence of class imbalance, which is a prevalent issue in prognostic datasets. This ensures a fair assessment of model predictions across both majority and minority classes. An Additional evaluation metric, Area Under the Precision-Recall Curve (AUPRC), is employed to provide a more complete understanding of the model behavior.
@@ 3.1 Datasets
AIforCOVID [65] Severity Outcome ($\mathbf{AFC}$) Mild: 839
@@
COVID-19-AR [66] ICU Admission ($\mathbf{CAR}$) Yes: 28
@@
CoCross [67] ICU Outcome ($\mathbf{CC}$) Alive: 243
@@
Stony Brook COVID-19 [68] Mortality ($\mathbf{NY_{small}}$) Alive: 1183
@@
In this work, we evaluate four publicly available COVID-19 chest X-ray (CXR) datasets selected for their relevance to prognosis prediction and their diversity in task definition, class imbalance, and sample size. Focusing on a common pathological context allows us to benchmark out selected set of model under consistent disease representation while capturing real-world challenges such as extreme data scarcity, heterogeneous imaging frequency, and clinical outcome variability.
Each dataset is annotated with different prognostic task, from mortality, severity, to ICU admission, and collectively spans balanced to highly imbalanced distributions and small to large sample sizes. Standardized care protocols during the pandemic ensure outcome consistency across datasets, enabling reliable comparisons [63, 64].
Table 1 reports the main statistics, including image count, patient numbers, and class distributions.
The AIforCOVID dataset [65] comprises 1,585 patients from six Italian hospitals, with two tasks: severity ($\mathbf{AFC}$), showing a balanced class distribution (53–47%), and mortality ($\mathbf{AFC_{m}}$), which is notably imbalanced (85–15%). Both tasks use the same patient cohort and CXRs.
The CoCross dataset [67] includes 387 CXRs from ICU patients monitored over time, with a binary ICU outcome label ($\mathbf{CC}$) that is moderately imbalanced (63–37%). The COVID-19-AR dataset [66], while limited to 99 CXRs, focuses on rural and underrepresented populations. Its ICU admission label ($\mathbf{CAR}$) presents a strong imbalance (71–29%), making it a valuable case for evaluating models in extreme low-data regimes.
Lastly, the Stony Brook COVID-19 dataset [68] is used for mortality prediction with two versions: $\mathbf{NY_{small}}$, comprising one CXR per patient closest to admission (1,365 CXRs, 87–13% imbalance), and $\mathbf{NY_{all}}$, containing all CXRs recorded during hospitalization (13,639 images, 64–36% imbalance). Both versions share the same patient population, allowing controlled comparisons of dataset size and class imbalance effects on prognostic modeling.
The selected datasets provide a suitable ground for prognostic assessment, incorporating key aspects frequently observed in prognosis tasks, such as high task complexity, variability in sample sizes, and different degrees of class imbalance. The influence of these prognostic data characteristics will be systematically analyzed and discussed in relation to the performance achieved by the various FMs across the different tasks, with the corresponding results presented in Section 4.
@@ 3.2 Models
For this study, we include a diverse set of models that have been pretrained using three distinct learning paradigms: supervised pretraining, self-supervised pretraining, and contrastive language-image pretraining. This design supports a controlled comparison of how architectural choices and pretraining strategies affect the transfer to prognosis tasks. A summary of the included models is provided in Table 2, alongside with further details about architectural class, pretraining strategy, pretraining dataset, vision backbone type, number of vision encoder parameters and embedding dimension, where the symbol “--” in the embedding dimension column indicates that the model is not Transformer-based or does not have a uniform embedding size.
@@
ResNet18 [69] Supervised ImageNet [14] ResNet18 11.7 - CNNs
ResNet50 [69] Supervised ImageNet [14] ResNet50 23.5 - CNNs
DenseNet121 [70] Supervised ImageNet [14] DenseNet121 7.9 - CNNs
DINOv2s [71] Self-supervised LVD-$142$[71] ViT-S/14 21 384 FMs
DINOv2b [71] Self-supervised LVD-$142$[71] ViT-B/14 86 512 FMs
DINOv2l [71] Self-supervised LVD-$142$[71] ViT-L/14 300 768 FMs
CLIP-Large [21] Contrastive LAION-400M [72] ViT-L/14 300 768 FMs
MedCLIPc [39] Contrastive CheXpert + MIMIC [73, 41] ResNet50 23 - FMs
MedCLIPv [39] Contrastive CheXpert + MIMIC [73, 41] Swin-T 27 - FMs
PubMedCLIP [74] Contrastive ROCO  [75] ViT-B/14 86 512 FMs
BioMedCLIP [40] Contrastive PMC-15M [40] ViT-B/14 86 512 FMs
@@
We include both ResNet18 and ResNet50 as representative CNN architectures due to their established success in transfer learning for medical imaging, particularly when pretrained on natural image datasets like ImageNet [15, 16]. Their low parameter count, strong inductive biases, and efficient training dynamics make them particularly effective in low-data settings, where overfitting and computational constraints are critical concerns. For the self-supervised learning (SSL) paradigm, we adopt DINOv2 [71], a state-of-the-art vision transformer (ViT)-based model that has shown competitive performance in medical imaging tasks [76, 55]. Its ability to extract rich, transferable features from unlabeled data makes it highly suitable for prognostic tasks in resource-constrained environments, where annotated datasets are often scarce. We include CLIP-Large [21] to assess the transferability of general-purpose contrastive language-image pretraining in prognosis prediction. Moreover, given the widespread adoption of contrastive pretraining also in the medical applications, we evaluate three medical pretrained variants: MedCLIP [39], PubMedCLIP [74], and BioMedCLIP [40] all pretrained on large-scale multimodal biomedical datasets. These models offer an ideal test ground to analyze domain-specific pretraining on our prognosis downstream tasks. For these multimodal architectures, we utilize only the vision encoder, integrating it with a task-specific linear adapter layer designed for each prognostic prediction task. This diverse selection, spanning architectural types, pretraining strategies, and domain alignments, enables a systematic evaluation of adaptation strategies and model robustness in extreme low-data prognostic settings.
@@ 3.3 PEFT, Linear Probing and Full Fine-Tuning
In the context of medical imaging prognosis, adapting deep learning models to clinical outcome prediction remains a challenging task that requires balancing computational efficiency with effective knowledge transfer. Given the characteristics of our datasets—limited sample size and class imbalance—we evaluate several PEFT techniques against standard LP and FFT, applied across models pretrained with diverse learning paradigms and datasets (Section 3.2). Table 3 summarizes the compatibility of each PEFT method with the considered architectures.
Low-Rank Adaptation (LoRA[32] introduces trainable low-rank matrices while keeping the pretrained weights frozen. By decomposing weight updates into two smaller matrices, it significantly reduces the number of trainable parameters while preserving the performance of the original model. This is particularly beneficial in prognosis tasks, where data scarcity makes overfitting a risk. Although commonly applied to transformers [77, 29], we also extend its evaluation to CNNs to assess its versatility across architectural types.
Vector-based Random Matrix Adaptation (VeRA[53] builds upon LoRA by employing fixed random projections shared across layers, modulated by trainable scaling vectors. This further reduces the parameter footprint while retaining competitive performance. Due to its reliance on a uniform feature space, VeRA is only applied to ViT-based models, excluding CNNs and the hierarchical Swin-based MedCLIPv.
BitFit [31] is a sparse fine-tuning method that updates only the bias terms of the model. Despite modifying fewer than 0.1% of the parameters, it often yields strong results, making it particularly attractive for prognosis applications where training resources and data availability are limited. Moreover, it offers insights into how far inductive biases encoded in pretrained models can be selectively activated through minimal adaptation.
Infused Adapter by Inhibiting and Amplifying Inner Activations (IA3[33] adapts a pretrained model by inserting multiplicative gates that modulate intermediate activations. These scaling vectors act on attention and feed-forward layers without altering the model’s main parameters. Initialized to preserve the model’s original behavior, IA3 enables stable and efficient adaptation, especially effective in low-data scenarios typical of prognosis tasks.
In addition to PEFT, we include FFT and LP as reference fine-tuning settings. FFT updates all parameters and serves as the upper-bound useful to evaluate the full model adaptability to new downstream tasks, while LP freezes the backbone and optimizes only a linear classification head, offering a lower-bound that evaluates the quality of pretrained features [21, 78]. Using both benchmarks establishes a performance spectrum, within which we assess the efficiency and expressiveness of PEFT strategies in real-world clinical settings.
@@ 3.4 Few-Shot Learning Analysis
FSL refers the problem setting in which a model must learn to generalize from only a few labeled examples per class [79]. This paradigm is particularly relevant in medical imaging, where annotated data is often scarce due to high labeling costs and limited expert availability. To evaluate the robustness and adaptability of PEFT methods under such constraints, we design a systematic few-shot learning benchmark. Let $\mathcal{Y}=\{1,\dots,C\}$ denote the set of class labels. For a fixed number of shots $k$, the few-shot training set is defined as follows:
@@
where exactly $k\in\mathbb{N}$, represents the samples drawn for each class, leading to a total of $kC$ training examples per experiment. We evaluate all PEFT strategies listed in Table 3, excluding FFT due to its tendency to overfit in low-data regimes. The training protocol follows a geometric progression over the number of shots: $k\in\{2,4,8,16,32\}$. This allows us to systematically characterize the scaling behavior of each method as the number of labeled samples increases. By exploring this setup, we aim to determine the extent to which PEFT approaches can effectively adapt to prognosis prediction tasks with extremely limited data, providing insights into their feasibility for real-world medical applications where annotated data could be extremely scarce.
@@ 3.5 Experimental Setup
All experiments followed a uniform configuration for the fine-tuning. Optimization was performed using Stochastic Gradient Descent (SGD) with a momentum of $0.9$, weight decay of $10^{-4}$, and an initial learning rate of $10^{-3}$. Training proceeded for a maximum of 1000 epochs, with a 50-epoch warm-up phase, and employed early stopping with a patience of 50 epochs to mitigate overfitting. We used a batch size of 128 for larger tasks (AFC, AFCm, NYsmall, and NYall), and 32 for smaller datasets (CAR, CC) to match their lower dimensionality.
@@
This variation helps assess the sensitivity of each method respect dimensionality of the adaptation space and how it is impacting the adaptation on the downstream tasks performance. All models employed in this study were equipped with a task-specific classification head, appended to the output of the vision backbone. Feature representations were extracted from the final block of each model and passed through a lightweight linear layer to produce the final logits. In the case of Vision Transformer (ViT) and foundation model (FM) architectures, we consistently used the class token as the global feature descriptor. This design choice ensures structural and dimensional homogeneity across all architectures, facilitating fair comparisons across PEFT strategies and model types.
@@
Regarding evaluation metrics, we adopted the Matthews Correlation Coefficient (MCC) as our primary metric due to its robustness in extreme class imbalance conditions [80]. The metric ranges from -$\text{-}1$to $1$, where $1$ indicates perfect prediction, 00 corresponds to random guessing, and -$\text{-}1$ denotes total disagreement between predictions and ground truth. Indeed, it considers true and false positives and negatives simultaneously, being well-suited for prognosis tasks where the minority class is typically of greater clinical interest yet severely underrepresented in the data (e.g., progression or adverse outcome). To complement MCC, we also report the Precision-Recall Area Under the Curve (PR-AUC), which is more sensitive than ROC-AUC to the performance on the positive (minority) class and is thus better aligned with the requirements of our benchmark’s tasks [81]. From the hardware perspective, all experiments were conducted using the PyTorch framework on a cluster equipped with 10 NVIDIA TESLA A40 GPUs. The complete code-base is publicly available at https://github.com/fruffini/PEFT_Prognosis.

### v2

@@ 3 Materials & Methods
We employ four publicly available COVID-19 CXR datasets annotated with a variety of prognostic outcomes, including mortality, severity, and ICU admission. These datasets cover a wide range of clinical scenarios and data challenges, including varying degrees of class imbalance, outcome heterogeneity, and differences in sample size, ranging from highly underrepresented cases to larger, more balanced datasets. This diversity reflects the complexities typical in prognosis prediction and enables a robust evaluation of model adaptability under different data regimes [mamandipoor2022covid19, jiao2021prognosticationCOVID19]. Further details of the datasets are provided in Section 3.1.
@@
We systematically evaluate models grouped into two main architectural families. The first group composed of three CNNs pretrained on ImageNet using supervised learning, as a strong baselines in medical imaging. The second group composed entirely by FMs, pretrained through self-supervised or contrastive learning, either on large-scale general image-text datasets or specialized medical datasets. These two group allow us to explore the interplay between architectural design, pretraining paradigms, and domain specificity in prognosis prediction tasks. Further details of the models are provided in Section 3.2.
@@
Each model undergoes adaptation to the target prognostic tasks using one of several fine-tuning strategies. These include FFT as a performance upper bound; LP to evaluate the quality of frozen representations; and multiple PEFT techniques, which are designed to minimize computational overhead and improve adaptation efficiency. The full range of supported configurations and their applicability to specific architectures are detailed in Section 3.
@@
To assess the generalization capabilities of models under realistic constraints, we benchmark each configuration under two training regimes. In the full-data regime, where models are trained using all available labeled samples. And the FSL regime, where we limit the training to few samples per class, simulating extreme data scarcity common in rare disease contexts or emerging health crises. This dual evaluation framework enables a comprehensive analysis of model robustness and the effectiveness of fine-tuning strategies across a spectrum of data availability scenarios.
@@
We evaluate model performance primarily using the Matthews Correlation Coefficient (MCC), a metric known for its robustness in the presence of class imbalance, which is a prevalent issue in prognostic datasets. This ensures a fair assessment of model predictions across both majority and minority classes. An Additional evaluation metric, Area Under the Precision-Recall Curve (AUC-PR), is employed to provide a broader understanding of the models’ behavior.
@@ 3.1 Datasets
AIforCOVID [soda2021aiforcovid-afc] Severity Outcome ($\mathbf{AFC}$) Mild: 839
@@
COVID-19-AR [desai2020chest-car] ICU Admission ($\mathbf{CAR}$) Yes: 28
@@
CoCross [kilintzis2022cocross-cc] ICU Outcome ($\mathbf{CC}$) Alive: 243
@@
Stony Brook COVID-19 [saltz2021stony-ny] Mortality ($\mathbf{NY_{small}}$) Alive: 1183
@@
In this study, we evaluate four publicly available COVID-19 chest X-ray (CXR) datasets, chosen for their relevance to prognosis prediction and their diversity in task definition, class distribution, and sample size. We focus on COVID-19 due to its representativeness and the presence of standardized care protocols established during the pandemic, which ensure consistent outcome definitions across datasets and enable reliable cross-cohort comparisons [mamandipoor2022covid19, jiao2021prognosticationCOVID19]. This design allows us to benchmark the selected models within a unified disease framework while still reflecting real-world challenges such as data scarcity, heterogeneous imaging frequency, and variability in clinical outcomes.
Each dataset is annotated with different task, ranging from mortality or severity, to ICU admission, and collectively spanning from balanced to highly imbalanced class distributions and small to large sample sizes. Table 1 reports the main statistics, including image count, patient numbers, and class distributions of the selected datasets and related tasks.
The AIforCOVID dataset [soda2021aiforcovid-afc] includes 1,585 patients collected from six Italian hospitals and defines two distinct prognostic tasks. The first concerns disease severity classification, Mild vs. Severe (AFC), featuring a balanced class distribution ($53–47\%$), while the second focuses on mortality prediction (AFCm), characterized by a pronounced class imbalance ($85–15\%$). Both tasks are derived from the same patient cohort and share identical CXRs examinations.
The CoCross dataset [kilintzis2022cocross-cc] comprises 387 CXRs from ICU patients monitored longitudinally, annotated with a binary ICU outcome label (CC) that exhibits moderate class imbalance ($63–37\%$). The COVID-19-AR dataset [desai2020chest-car], although limited to 99 CXRs, specifically targets rural and underrepresented populations. Its ICU admission label (CAR) is strongly imbalanced (71–29%), providing a valuable testbed for assessing model robustness under extreme low-data and class-imbalance conditions.
Lastly, the Stony Brook COVID-19 dataset [saltz2021stony-ny] is employed for mortality prediction and includes two complementary versions: $\mathbf{NY_{small}}$, consisting of one CXR per patient taken closest to hospital admission (1,365 CXRs; $87–13\%$ class imbalance), and $\mathbf{NY_{all}}$, comprising all CXRs acquired throughout hospitalization (13,639 images; $64–36\%$ class imbalance). Both versions share the same patient cohort, enabling controlled investigations of how dataset size and class imbalance affect prognostic model performance.
Collectively, these datasets offer a comprehensive testbed for prognostic evaluation, encompassing key characteristics typical of real-world prognosis tasks, namely, high task complexity, variability in sample size, and differing degrees of class imbalance. The influence of these factors on model behavior will be systematically examined and discussed in relation to the performance of the evaluated Foundation Models across all tasks, with detailed results reported in Section 4.
@@ 3.2 Models
For this study, we evaluate a diverse set of models pretrained under three distinct paradigms: supervised, self-supervised, and contrastive language-image learning. This setup enables a controlled comparison of how different architectural designs and pretraining strategies influence model transferability to prognostic tasks. A comprehensive summary of the evaluated models is provided in Table 2, detailing their architectural class, pretraining strategy, pretraining dataset, vision backbone, number of vision encoder parameters, and embedding dimension. The symbol - in the embedding dimension column indicates models that are either non Transformer-based or do not possess a uniform embedding size.
@@
ResNet18 [he2016deep-resnet] Supervised ImageNet [ImageNet] ResNet18 11.7 - CNNs
ResNet50 [he2016deep-resnet] Supervised ImageNet [ImageNet] ResNet50 23.5 - CNNs
DenseNet121 [huang2017densely-densenet] Supervised ImageNet [ImageNet] DenseNet121 7.9 - CNNs
DINOv2s [oquab2023dinov2] Self-supervised LVD-142142[oquab2023dinov2] ViT-S/14 21 384 FMs
DINOv2b [oquab2023dinov2] Self-supervised LVD-142142[oquab2023dinov2] ViT-B/14 86 512 FMs
DINOv2l [oquab2023dinov2] Self-supervised LVD-142142[oquab2023dinov2] ViT-L/14 300 768 FMs
CLIP-Large [radford2021learning-cliplarge] Contrastive LAION-400M ViT-L/14 300 768 FMs
MedCLIPc [wang2022medclip] Contrastive CheXpert + MIMIC [johnson2023mimic] ResNet50 23 - FMs
MedCLIPv [wang2022medclip] Contrastive CheXpert + MIMIC [johnson2023mimic] Swin-T 27 - FMs
PubMedCLIP [eslami-etal-2023-pubmedclip] Contrastive ROCO  [ruckert2024rocov2] ViT-B/14 86 512 FMs
BioMedCLIP [zhang2023biomedclip] Contrastive PMC-15M [zhang2023biomedclip] ViT-B/14 86 512 FMs
@@
We include both ResNet18 and ResNet50 as representative CNN architectures, given their well-established success in transfer learning for medical imaging, particularly when pretrained on large-scale natural image datasets such as ImageNet [ke2021chextransfer, raghu2019transfusion_review]. Their relatively low parameter count, strong inductive biases, and efficient training dynamics make them especially suitable for low-data scenarios, where overfitting and computational constraints are major challenges.
For the self-supervised learning (SSL) paradigm, we adopt DINOv2 [oquab2023dinov2], a state-of-the-art Vision Transformer (ViT) model that has demonstrated competitive performance in medical imaging applications [huang2024comparative, jin2024fairmedfm]. Its capacity to learn rich and transferable visual representations from unlabeled data makes it particularly valuable for prognostic tasks in resource-limited settings, where annotated data are often scarce. Lastly, we include CLIP-Large [radford2021learning-cliplarge] to evaluate the generalization potential of large-scale contrastive language–image pretraining for prognosis prediction. For the contrastive pretraining, we selected also a set of models pretrained on large-scale multimodal biomedical datasets: MedCLIP [wang2022medclip], PubMedCLIP [eslami-etal-2023-pubmedclip], and BioMedCLIP [zhang2023biomedclip].
For the multimodal architectures (i.e., MedCLIP), we extracted the vision encoder from the whole architecture, and later adding a task-specific linear adapter layer designed for each prognostic prediction task. This diverse selection, spanning architectural types, pretraining strategies, and domain alignments, enables a systematic evaluation of adaptation strategies and model robustness in extreme low-data prognostic settings.
@@ 3.3 PEFT, Linear Probing and Full Fine-Tuning
In the context of medical imaging prognosis, adapting deep learning models to clinical outcome prediction remains a challenging task that requires balancing computational efficiency with effective knowledge transfer. Given the characteristics of our datasets, we evaluate several PEFT techniques against standard LP and FFT, across models pretrained with diverse learning paradigms and datasets (Section 3.2). Table 3 summarizes the compatibility of each PEFT method with the considered architectures. Detailed mathematical formulations of all methods are provided in Appendix A.
Low-Rank Adaptation (LoRA[hu2021lora] introduces trainable low-rank matrices while keeping the pretrained weights frozen. By decomposing weight updates into two smaller matrices, it significantly reduces the number of trainable parameters while preserving the performance of the original model. This is particularly beneficial in prognosis tasks, where data scarcity makes overfitting a risk. Although commonly applied to transformers [he2021towards, han2024parameter], we also extend its evaluation to CNNs to assess its versatility across architectural types.
Vector-based Random Matrix Adaptation (VeRA[kopiczko2023vera] builds upon LoRA by employing fixed random projections shared across layers, modulated by trainable scaling vectors. This further reduces the parameter footprint while retaining competitive performance. Due to its reliance on a uniform feature space, VeRA is only applied to ViT-based models, excluding CNNs and the hierarchical Swin-based MedCLIPv.
BitFit [zaken2021bitfit] is a sparse fine-tuning method that updates only the bias terms of the model. Despite modifying fewer than $0.1\%$ of the parameters, it often yields strong results, making it particularly attractive for prognosis applications where training resources and data availability are limited. Moreover, it offers insights into how far inductive biases encoded in pretrained models can be selectively activated through minimal adaptation.
Infused Adapter by Inhibiting and Amplifying Inner Activations (IA3[liu2022-ia3] adapts a pretrained model by inserting multiplicative gates that modulate intermediate activations. These scaling vectors act on attention and feed-forward layers without altering the model’s main parameters. Initialized to preserve the model’s original behavior, IA3 enables stable and efficient adaptation, especially effective in low-data scenarios typical of prognosis tasks.
In addition to PEFT techniques, we include FFT and LP as reference fine-tuning settings. FFT updates all parameters during training, and serves as the upper-bound useful to evaluate how full model adaptability perform to prognosis downstream tasks, while LP freezes the backbone and optimizes only a linear classification head, offering a lower-bound that evaluates the quality of pretrained features [radford2021learning-cliplarge, he2020momentum]. Using both benchmarks establishes a performance spectrum, within which we assess the efficiency and expressiveness of PEFT strategies in real-world clinical settings.
@@ 3.4 Few-Shot Learning Analysis
FSL refers the problem setting in which a model must learn to generalize from only a few labeled examples per class [pachetti2024systematic]. This paradigm is particularly relevant in medical imaging, where annotated data is often scarce due to high labeling costs and limited expert availability. To evaluate the robustness and adaptability of PEFT methods under such constraints, we design a systematic few-shot learning benchmark. Let $\mathcal{Y}=\{1,\dots,C\}$ denote the set of class labels. For a fixed number of shots $k$, the few-shot training set is defined as follows:
@@
where exactly $k\in\mathbb{N}$, represents the samples drawn for each class, leading to a total of $kC$ training examples per experiment. We evaluate all PEFT strategies listed in Table 3, excluding FFT due to its tendency to overfit in low-data regimes. The training protocol follows a geometric progression over the number of shots: $k\in\{2,4,8,16,32\}$. This allows us to systematically characterize the scaling behavior of each method as the number of labeled samples increases. Importantly, while training sets are artificially balanced, test sets preserve the natural class imbalance of each dataset. This imbalance poses a realistic challenge: models learn under balanced supervision but must generalize to clinically imbalanced conditions. By exploring this setup, we aim to determine the extent to which PEFT approaches can effectively adapt to prognosis prediction tasks with extremely limited data.
@@ 3.5 Experimental Setup
All experiments followed a uniform configuration for the fine-tuning. Optimization was performed using Stochastic Gradient Descent (SGD) with a momentum of $0.9$, weight decay of $10^{-4}$, and an initial learning rate of $10^{-3}$. Training proceeded for a maximum of 1000 epochs, with a 50-epoch warm-up phase, and employed early stopping with a patience of 50 epochs to mitigate overfitting. We used a batch size of 128 for larger tasks (AFC, AFCm, NYsmall, and NYall), and 32 for smaller datasets (CAR, CC) to match their lower dimensionality. To ensure strict experimental determinism and reproducibility, random seeds were fixed for all stages of experimentation, including data splitting, model initialization, and data loading/batching.
@@
This variation helps assess the sensitivity of each method respect dimensionality of the adaptation space and how it is impacting the adaptation on the downstream tasks performance.
All models employed in this study were equipped with a task-specific classification head, appended to the output of the vision backbone. Feature representations were extracted from the final block of each model and passed through a lightweight linear layer to produce the final logits. In the case of Vision Transformer (ViT) and foundation model (FM) architectures, we consistently used the class token as the global feature descriptor. This design choice ensures structural and dimensional homogeneity across all architectures, facilitating fair comparisons across PEFT strategies and model types.
@@
Regarding evaluation metrics, we adopted the Matthews Correlation Coefficient (MCC) as our primary metric due to its robustness in extreme class imbalance conditions [chicco2023matthews]. The metric ranges from -$\text{-}1$to $1$, where $1$ indicates perfect prediction, 0 corresponds to random guessing, and -$\text{-}1$ denotes total disagreement between predictions and ground truth. Indeed, it considers true and false positives and negatives simultaneously, being well-suited for prognosis tasks where the minority class is typically of greater clinical interest yet severely underrepresented in the data (e.g., progression or adverse outcome). To complement MCC, we also report the Precision-Recall Area Under the Curve (PR-AUC), which is more sensitive than ROC-AUC to the performance on the positive (minority) class and is thus better aligned with the requirements of our benchmark’s tasks [PRAUC-1].
Experiments were conducted on a high-performance computing cluster equipped with 10 NVIDIA TESLA A40 GPUs (48GB each). The total compute required for the full benchmark was approximately 3,200 GPU-hours, corresponding to an approximated average of 6 GPU-hours per model-dataset combination. For all the models presented in this benchmark, we used publicly released pretrained weights available on their official repositories or Hugging Face hub. The corresponding links to the official repositories are reported in the appendix Table 8. The complete code-base is publicly available at https://github.com/fruffini/PEFT_Prognosis.